# DiffGAN: A Test Generation Approach for Differential Testing of Deep Neural Networks

### v2

@@ 7 Threats to Validity
External threats to validity concern the generalizability of our study. We mitigate this threat by considering eight different combinations of widely used models and datasets. These combinations feature diverse model pairs varying in architecture, training data, optimizers, and hyperparameters. Additionally, we explored a wide range of testing budgets (execution times) in our experiments and compared our results against a SOTA baseline for the differential testing of DNN models.
@@ 9 Conclusion
For future work, we aim to enhance the generalizability of DiffGAN by extending it to support a wider range of models, machine learning tasks, and input types. We also aim to explore DiffGAN in dynamic environments to assess the side effects of updating a single ML component on the overall system performance. Rather than comparing pairs of DNN models, we will compare the original system with the updated version. This approach will provide valuable insights into maintaining system reliability and adaptability after updates, particularly in critical areas like autonomous driving, and healthcare.

### v3

@@ 7 Threats to Validity
External threats to validity concern the generalizability of our study. We considered in our experiments two datasets and eight pairs of DNN models, which may limit the generalizability of our study. While we recognize the potential of adding more datasets, such as ImageNet or Tiny ImageNet, incorporating them would require a substantial increase in manual effort to validate the generated images, which was already extensive. For instance, Tiny ImageNet involves 200 classes, making manual validation extremely challenging. Accurately classifying each image among so many visually similar categories is not only time-consuming but also difficult to perform reliably without expert annotators. This is precisely why we chose datasets where such validation, though still effort-intensive, remained feasible. To further mitigate the generalizability threat of our study, we included in our experiments 16 models across diverse datasets (MNIST and CIFAR-10), multiple domains, and various architectures (e.g., LeNet, ResNet, Inception), all of which are widely used in the literature. Furthermore, these combinations of models feature diverse model pairs varying in training data, optimizers, and hyperparameters. Our evaluation includes both highly accurate model pairs and relatively less accurate ones to enhance the generalizability of our results. Additionally, we explored a wide range of testing budgets (execution times) in our experiments and compared our results against a SOTA baseline for the differential testing of DNN models.
@@ 8 Related Work
Several generative model-based approaches for test generation of DNNs have been proposed in the literature [76, 19, 77]. For instance, Dola et al. [76] introduced a black-box GAN-based method aimed at generating test inputs that are both feature-diverse and include rare cases. Their approach leverages a generative model and applies combinatorial interaction testing [78] to the model’s latent space to enhance the diversity of the generated test inputs. Also, Zhang et al. [19] proposed DeepRoad, a GAN-based approach for testing DNN-based autonomous driving systems (ADS). DeepRoad generates driving scenes with different weather conditions and applies metamorphic testing to detect mispredictions. Metamorphic relations are defined such that driving behavior remains consistent, regardless of how scenes are altered to simulate different weather conditions. Similarly, Li et al. [77] proposed TACTIC, a test generation approach for DNN-based ADS. TACTIC leverages a GAN-based image-to-image translation framework to identify critical weather conditions that increase the likelihood of erroneous system behavior. While these approaches leverage various types of GANs to generate test inputs for DNNs, they focus on identifying mispredictions in individual DNN models. In contrast, our approach is specifically designed for differential testing between pairs of DNN models. We generate triggering inputs that expose behavioral disagreements between models, rather than targeting mispredictions in a single model.
@@ 9 Conclusion
For future work, we aim to enhance the generalizability of DiffGAN by extending it to support a wider range of models, machine learning tasks, and input types. We also aim to explore DiffGAN in dynamic environments to assess the side effects of updating a single ML component on the overall system performance. Rather than comparing pairs of DNN models, we will compare the original system with the updated version. This approach will provide valuable insights into maintaining system reliability and adaptability after updates, particularly in critical areas like autonomous driving and healthcare.
@@ Acknowledgments
Appendix I
Diversity Comparison
We evaluated the input diversity generated by DFlare, DRFuzz, and our approach on the MNIST dataset using four model pairs. Following the methodology of RQ3, we employed Geometric Diversity and Shannon Entropy as metrics, with a fixed generation time of one hour for each method. As shown in Table XIII, the results indicate that DFlare produced highly redundant inputs compared to both DRFuzz and DiffGAN, consistently across all model pairs. Due to its poor performance in generating diverse inputs, we excluded DFlare from further analysis.
Table XIII: Diversity scores of MNIST (1h Configuration)
DFlare DRfuzz DiffGAN
Models Configs GD Exponential Shannon ($e^{H}$) GD Exponential Shannon ($e^{H}$) GD Exponential Shannon ($e^{H}$)
M1, M2 1h -377.52 $e^{4.74}\approx 114.88$ 422.03 $e^{11.36}\approx 86,361.86$ 530.62 $e^{12.10}\approx 178,632.90$
M3, M4 1h 132.15 $e^{4.75}\approx 115.07$ 438.42 $e^{11.39}\approx 89,882.98$ 727.20 $e^{12.21}\approx 200,195.46$
M5, M6 1h -72.37 $e^{4.77}\approx 117.88$ 410.67 $e^{11.50}\approx 98,610.81$ 712.76 $e^{12.27}\approx 213,402.74$
M7, M8 1h 36.42 $e^{4.79}\approx 120.37$ 116.72 $e^{11.40}\approx 90,159.89$ 481.94 $e^{12.18}\approx 178,751.64$